You are flipping a coin 3 times and rolling a 6-sided dice once. How many different possible outcomes are there?
Answer by greenestamps(13198)



The experiment is not described clearly enough.

There are obviously 6 possible outcomes on rolling a 6-sided die once.

For flipping a coin 3 times, the number of possible "outcomes" depends on whether it means the number of possible SEQUENCES of heads/tails (2*2*2=8 "outcomes"), or whether it means the NUMBERS of heads and tails (0 or 1 or 2 or 3 --> 4 "outcomes").

So I see two possible answers to the problem:

sequences of heads/tails and number on the die: 8*6=48 different possible outcomes
numbers of heads/tails and number on the dies: 4*6=24 different possible outcomes